OpenStudy (anonymous):

Find the integral [(3x-2)^4]/[7]

OpenStudy (anonymous):

Is there an easier way to expand it?

OpenStudy (anonymous):

\[\frac{ (3x-2)^{4} }{ 7 } dx\]

OpenStudy (anonymous):

yeah try \(u=3x-2,du=3dx,\frac{1}{3}u=dx\)

OpenStudy (anonymous):

you get \[\frac{1}{21}\int u^4du\] and

OpenStudy (anonymous):

Ah! yes okay thanks!!!!!!

OpenStudy (anonymous):

so would it be \[\frac{ (3x-2)^{5} }{ 135 } +c\]

OpenStudy (anonymous):

i think your denominator is wrong

OpenStudy (anonymous):

\(21\times 5=105\)

OpenStudy (anonymous):

oh haha I wrote down 27 for some reason

OpenStudy (anonymous):

∫(3x-2)^4/[7] dx 1/7∫(3x-2)^5dx so, 1/7{(3x-2)^5/(5*3)}+c therefore dinominator '3' is derivative of base. finally; 1/135(3x-2)^5 +c .